Consider the equation x^2-18x=3. What number must be added to both sides of the equation in order to complete the square.
The answer is one of the following. Which one is it?
A) 9
B) 18
C) 27
D) 81

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52780)

.

81.    ANSWER


CHECK.    x^2 - 18x + 81 = (x-9)^2.


Answer by Alan3354(69443)

c = (b/2a)^2 = (-18/2)^2 = 81